OpenStudy (anonymous):

a typical pyrex glass contains 12.9% of b2o3 and 80.7% of sio2 (by weight) besides other oxides (al2o3 , na2O , k2O) what is the weight ratio of silicon to boron in the glass ( atomic weight si=28.1 b=10.8)?

OpenStudy (anonymous):

options are 9.32:1 9.76:1 9.29:1 9.43:1

OpenStudy (anonymous):

the answer is 9.43:1 but i dont know to solve it

OpenStudy (anonymous):

Empirical Formula question \[B_{2}O_{3} : SiO_{2}\] \[69.7 : 60.1\] ( find the molecular mass of each....) and then find their percentage like for example ratio of B203 \[\frac{12.9}{100} \times 69.7= 9\] then for SiO2 \[\frac{80.3}{100} \times 60= approximately.. u ..get.. 48..but .\] \[according.2..ur..answer..u.. can..assume ..\it..as..43:)\]

OpenStudy (anonymous):

AND forgot to mention....the other compounds they dont make a big issue (in the multiple choic...coz all of them have the number "1"

OpenStudy (anonymous):

take a 100 g sample...12.9 g of b2o3 will be there...find the mass of b in it....similarly find for Si and thus find the ratio..

OpenStudy (anonymous):

can u solve the mass for b plzz just an example

OpenStudy (anonymous):

 n gives the no of moles of b2o3...multiply it by molar mass of B times 2 to bet mass of total B atoms in terms of X which is approx X/70 (i took x g sample instead of 100g) ....similarly do for Si and cancel out x in thee ratio...